OpenStudy (anonymous):

The table shows data gathered by an environmental agency about the decreasing depth of a lake during the first few months of a drought. Time (weeks) Lake Depth (feet) 0 346.0 2 344.8 5 343.0 7 341.8 10 340.0 12 338.8 The lake's depth after four weeks will be

OpenStudy (anonymous):

You need to graph this

OpenStudy (anonymous):

Okay, done it, here is the result: just entered it into: http://www.alcula.com/calculators/statistics/scatter-plot/

OpenStudy (anonymous):

This shows that there is a linear correlation between each of the points

OpenStudy (anonymous):

Now wait, I can get the answer

OpenStudy (anonymous):

I already got it, but thank you

OpenStudy (anonymous):

here is new plot with equation:

OpenStudy (anonymous):

So our equation is y= 346 - 0.6x i.e depth = 346 -0.6 (time in weeks) depth = 346 -0.6 (4 weeks) i.e. we lookin for depth after 4 weeks. depth(y) = 346 - 0.6(4) y= 346 - 2.40 346-2.40 = 343.6